# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

### v3

@@ C.1. Additional Accuracy and Fairness Metrics of Predictive Model
In this section, we extend the results discussed in Section 4 by presenting plots for additional popular accuracy and fairness metrics of downstream model performance. Figures 18, 19, 21, and 20 illustrate Accuracy, Disparate Impact, True Negative Rate Difference, and Selection Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms. Fairness metrics are calculated for an intersectional group when multiple sensitive attributes are present in the dataset, and for a single group when there is only one sensitive attribute (see Table 2 for reference).
@@
Figure 19. Disparate Impact (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.
@@
Figure 20. Selection Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.
@@
Figure 21. True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.
@@ D.1. Fixed Train and Test Missingness Fractions
Fairness of predictive model. The fairness of predictive models according to true positive rate difference is illustrated in Figure 26. The plot reveals that model fairness is highly sensitive to missingness shift. For instance, for the diabetes dataset, a predictive model trained on an MCAR train set can exhibit worse TPRD on an MNAR test set compared to an MCAR test set. Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set. We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 27-44. These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across all ML and DL-based imputers and all our datasets.
@@ D.2. Variable Train and Test Missingness Rates
Other fairness metrics under variable train set missingness. Figures 48, 49, and 50 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
@@
Other fairness metrics under variable test set missingness. Figures 51, 52, and 53 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.
The plots reinforce the assertion in Section 5.2 that model fairness is highly sensitive to missingness shift. For example, the subfigures for MCAR train & MNAR test and MAR train & MNAR test show the largest deviations with increasing test error rates for all three fairness metrics, while plots for other settings remain relatively flat. Interestingly, when null imputers and models are fitted on the MNAR train set, model performance on the MNAR test set across different test error rates do not change significantly compared to MCAR train & MNAR test and MAR train & MNAR test settings. This indicates that the missingness mechanism in the train set can significantly affect model fairness on the MNAR test set. Additionally, miss-forest and auto-ml proved to be the most robust MVI techniques against test set missingness shifts under all conditions.

### v4

@@ C.2. Additional Accuracy and Fairness Metrics of Predictive Model
In this section, we extend the results discussed in Section 4 by presenting plots for additional popular accuracy and fairness metrics of downstream model performance. Figures 20, 21, 22, and 23 illustrate Accuracy, Disparate Impact, Selection Rate Difference, and True Negative Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms. Figures 24 and 25 present scatter plots for all imputers that extend scatter plots from Section 4.4. Fairness metrics are calculated for an intersectional group when multiple sensitive attributes are present in the dataset, and for a single group when there is only one sensitive attribute (see Table 2 for reference).
@@
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 21. Disparate Impact (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are ideal/fair. Datasets are ordered in increasing order by size. The blue line indicates the performance of a model trained on clean data.
@@
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 22. Selection Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The blue line indicates the performance of a model trained on clean data.
@@
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 23. True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots).
(a) F1 (imputation) vs F1 (model)
(b) RMSE (imputation) vs F1 (model)
(c) KL divergence (imputation) vs F1 (model)
(a) F1 score difference (imputation) vs TPRD (model)
(b) RMSE difference (imputation) vs TPRD (model)
(c) KL divergence difference (imputation) vs TPRD (model)
@@ D.1. Fixed Train and Test Missingness Fractions
Fairness of predictive model. The fairness of predictive models according to true positive rate difference is illustrated in Figure 30. The plot reveals that model fairness is highly sensitive to missingness shift. For instance, for the diabetes dataset, a predictive model trained on an MCAR train set can exhibit worse TPRD on an MNAR test set compared to an MCAR test set. Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set. We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 31-48. These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across ML- and DL-based imputers and all our datasets.
@@ D.2. Variable Train and Test Missingness Rates
Other fairness metrics under variable train set missingness. Figures 52, 53, and 54 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
@@
Other fairness metrics under variable test set missingness. Figures 55, 56, and 57 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.
The plots reinforce the assertion in Section 5.2 that model fairness is highly sensitive to missingness shift. For example, the subfigures for MCAR train & MNAR test and MAR train & MNAR test show the largest deviations with increasing test error rates for all three fairness metrics, while plots for other settings remain relatively flat. Interestingly, when null imputers and models are fitted on the MNAR train set, model performance on the MNAR test set across different test error rates do not change significantly compared to MCAR train & MNAR test and MAR train & MNAR test settings. This indicates that the missingness mechanism in the train set can significantly affect model fairness on the MNAR test set. Additionally, miss-forest, auto-ml, and nomi proved to be the most robust MVI techniques against test set missingness shifts under all conditions.